OpenStudy (anonymous):

How would I write and expression for the following number sequence 4,6,10,18,34. I know that I am doubling number plus 2 that I had to each number.

OpenStudy (amistre64):

write it out as an equation, right or wrong, give it your best shot... then we can have something to work with

OpenStudy (amistre64):

i see +2 +4 +8 +6 even number right?

OpenStudy (amistre64):

4+2 +4 4+6 +8 4+14 +16 4+30 see a pattern?

OpenStudy (anonymous):

four plus 2 = 6 than six plus 2 square yes

OpenStudy (anonymous):

n squared plus 2?

OpenStudy (amistre64):

4+2 +4 4+6 +8 4+14 +16 4+30 ----------------- +32 4+62 would be next right?

OpenStudy (amistre64):

2^n is the right most column

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

2n + 2^n maybe?

OpenStudy (amistre64):

i was never good at these :)

OpenStudy (anonymous):

what do you mean? I have the pattern but the writing the expression is messing me up.

OpenStudy (amistre64):

the equation is just the pattern written out... really

OpenStudy (amistre64):

does n have to start at zero?

OpenStudy (amistre64):

" I know that I am doubling number plus 2 that I had to each number." is this your patter?

OpenStudy (amistre64):

2n is doubling a number... 2n + 2 is what I think you wrote....

OpenStudy (amistre64):

4,6,10,18,34 is this our only concern? its finite?

OpenStudy (anonymous):

ok try thus 2n square plus 2?

OpenStudy (amistre64):

2(1)^2 + 2 = 4 we missed "6" 2(2)^2 + 2 = 10

OpenStudy (amistre64):

2^n + ... 2^(1) + 2 = 4 2^(2) + 2 = 6 2^(3) + 2 = 10 2^(4) + 2 = 18 ....might be onto something here

OpenStudy (amistre64):

\[\sum_{1}^{4} (2^n +2)\]